OpenStudy (anonymous):

graph the inequality 2x-3y is less than or = to 3 .

OpenStudy (australopithecus):

http://openstudy.com/study#/updates/4f8ddb49e4b000310fac4cfc I explain how to graph these and find the inequalities here

OpenStudy (australopithecus):

or you can wait

OpenStudy (anonymous):

First step is to find the "equals" part. For two-variable linear inequalities, the "equals" part is the graph of the straight line. In this case, that means the "equals" part is the line 2x-3y = 3. Let's change the equation a bit so it's easier to graph shall we? 2x - 3y = 3 -3y = 3 - 2x y = (3/-3) - 2x/-3 y = (2/3)x - 1 Graphed, the "equals" part looks like  Now for the less than. What we want to show is that every y beneath that line is also a valid answer. To do that, we color it in! So the less than or equals looks kind of like: